JOINT INVERSE PROBLEMS AND DIFFERENTIAL GEOMETRY/PDE SEMINAR

Tuesday, July 20, 2004 Padelford C-36 1:30 pm

On a two dimensional non-linear Fourier transform

Alexandru Tamasan (University of Toronto, Canada)

I will discuss in detail the diagonalization of the Beals and Coifman formalism in inverse scattering. The scattering transform exhibits Fourier Transform like properties, thus justifying the title.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.